# Flexible and efficient spatial extremes emulation via variational autoencoders

### v3

@@ B.1 Properties of exponentially-tilted positive-stable variables
    $\log$ $F_{ij}\{q_{i}(t),q_{j}(t)\}=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k=1}^{K}\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}$
@@
    $=$ $\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k=1}^{K}\left[\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}t^{-1/\alpha}}{c^{\prime}_{i}{}^{1/\alpha}\left\{1+O(t^{1-1/\alpha})\right\}^{1/\alpha_{0}}}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}t^{-1/\alpha}}{c^{\prime}_{j}{}^{1/\alpha}\left\{1+O(t^{1-1/\alpha})\right\}^{1/\alpha_{0}}}\right]^{\alpha}$
    $=$ $\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k=1}^{K}\left[\gamma_{k}+\tau^{\frac{1}{\alpha_{0}}}\left(\frac{\omega_{ki}^{1/\alpha}}{c^{\prime}_{i}{}^{1/\alpha}}+\frac{\omega_{kj}^{1/\alpha}}{c^{\prime}_{j}{}^{1/\alpha}}\right)t^{-\frac{1}{\alpha}}+O\left(t^{1-\frac{2}{\alpha}}\right)\right]^{\alpha}$
    $=$ $\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k\in\bar{\mathcal{D}}}\left[\gamma_{k}^{\alpha}+\frac{\alpha\tau^{1/\alpha_{0}}}{\gamma_{k}^{1-\alpha}}\left(\frac{\omega_{ki}^{1/\alpha}}{c^{\prime}_{i}{}^{1/\alpha}}+\frac{\omega_{kj}^{1/\alpha}}{c^{\prime}_{j}{}^{1/\alpha}}\right)t^{-\frac{1}{\alpha}}+O\left(t^{1-\frac{2}{\alpha}}\right)\right]-$
    $\quad\tau^{\frac{\alpha}{\alpha_{0}}}\sum_{k\in\mathcal{D}}\left(\frac{\omega_{ki}^{1/\alpha}}{c^{\prime}_{i}{}^{1/\alpha}}+\frac{\omega_{kj}^{1/\alpha}}{c^{\prime}_{j}{}^{1/\alpha}}\right)^{\alpha}t^{-1}-O(t^{1-\frac{2}{\alpha}})$
@@
    $=$ $-d_{ij}t^{-1}-\left(\frac{c_{i}}{c^{\prime}_{i}{}^{1/\alpha}}+\frac{c_{j}}{c^{\prime}_{j}{}^{1/\alpha}}\right)t^{-\frac{1}{\alpha}}-O(t^{1-\frac{2}{\alpha}}),\;\text{as }t\rightarrow\infty.$
    It follows from (B.8) that
@@
    as $t\rightarrow\infty$, in which $d_{ij}=\tau^{{\alpha}/{\alpha_{0}}}\sum_{k\in\mathcal{D}}\{(\omega_{ki}/c^{\prime}_{i})^{1/\alpha}+(\omega_{kj}/c^{\prime}_{j})^{1/\alpha}\}^{\alpha}\in(1,2)$.
    If $\mathcal{C}_{i}\cap\mathcal{C}_{j}\neq\emptyset$, we know from (B.8) that $\chi_{ij}=2-d_{ij}\in(0,1)$ and
@@
is a limiting measure that occurs in the limiting distribution for normalized maxima. It is used to describe the multivariate extremal dependence of a spatial process, and the $n_{s}$-dimensional extremal coefficient $V(1,\ldots,1)$ is of particular interest. This extremal coefficient has a range of $[1,n_{s}]$, with the lower and upper ends indicating, respectively, perfect dependence and independence. As a polarized case, if $\gamma_{k}>0$, for all $k=1,\ldots,K$, then $\mathcal{C}_{j}\cap\mathcal{D}=\emptyset$ for all $j$’s, and thus we have
@@
Here, we can approximate $q_{j}(t)$ using the results from Corollary 1.1.1. From Proposition 1.1, we can deduce that $V(1,\ldots,1)=n_{s}$, which corresponds to joint extremal independence. By contrast, if all $\gamma_{k}=0$ and one knot covers the entire spatial domain, we have $V(1,\ldots,1)\in(1,n_{s})$, which corresponds to joint extremal dependence.
@@ C.1 Monte Carlo estimates of $\mathrm{ARE}_{\psi}(u)$
Proof of Theorem 2.1.
@@
From Expression (C.5), we see that the asymptotic normality of $n^{1/2}\{\widehat{\mathrm{ARE}}_{\psi}(u)-\mathrm{ARE}_{\psi}(u)\}$ is also guaranteed. However, the exact expression of its asymptotic variance requires a much more careful examination of the correlations among the ranks of $\boldsymbol{X}_{i}$, $i=0,1,\ldots,n_{g}$; that is, we need to device a multivariate linear rank statistics of $\boldsymbol{X}_{i}$, $i=0,1,\ldots,n_{g}$; see Ruymgaart and van Zuijlen (1978).
@@ C.2 Convergence of $\mathrm{ARE}_{\psi}(u)$
Proof of Theorem 2.2.
By the definition of the tail dependence measure in Eq. (9),

### v4

@@ A.1 Properties of exponentially-tilted positive-stable variables
    $\begin{split}\log&F_{ij}\{q_{i}(t),q_{j}(t)\}=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k=1}^{K}\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}\\ &=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k\in\bar{\mathcal{D}}}\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}-\sum_{k\in\mathcal{D}}\left\{\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}.\end{split}$ (A.14)
    For the second summation, the approximation in Eq. (A.12) still holds, except that the residual term becomes $o\left(t^{-\frac{2}{\alpha}}\right)$ due to the asymptotic approximations in Eq. (A.13). In the following, we examine the third summation by conditioning on whether $\mathcal{C}_{i}\cap\mathcal{C}_{j}\cap\mathcal{D}=\emptyset$ or $\mathcal{C}_{i}\cap\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$.
    If $\mathcal{C}_{i}\cap\mathcal{C}_{j}\cap\mathcal{D}=\emptyset$, then locations $i$ and $j$ are not covered by the same compact basis function with $\gamma_{k}=0$ even though $\mathcal{C}_{i}\cap\mathcal{D}\neq\emptyset$ and $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$ (i.e., they are individually impacted by different heavy-tailed expPS variables). In this case:
@@
    $\begin{split}\sum_{k\in\mathcal{D}}\left\{\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}&=\sum_{k\in\mathcal{D}}\frac{\tau^{\alpha/\alpha_{0}}\omega_{ki}}{q_{i}^{\alpha/\alpha_{0}}(t)}+\sum_{k\in\mathcal{D}}\frac{\tau^{\alpha/\alpha_{0}}\omega_{kj}}{q_{j}^{\alpha/\alpha_{0}}(t)}=c_{i}^{\prime}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{\alpha}{\alpha_{0}}}+c_{j}^{\prime}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{\alpha}{\alpha_{0}}}.\\ \end{split}$
    Combine this result with Eq. (A.12) and feed them in Eq. (A.14), we have
    11 11 $-F_{ij}\{q_{i}(t),q_{j}(t)\}=$
    $=1-\exp\left[c_{i}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{1}{\alpha_{0}}}+c_{j}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{1}{\alpha_{0}}}+c_{i}^{\prime}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{\alpha}{\alpha_{0}}}+c_{j}^{\prime}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{\alpha}{\alpha_{0}}}\right.$
    $\left.+d_{i}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{2}{\alpha_{0}}}+d_{j}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{2}{\alpha_{0}}}+c_{ij}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{1}{\alpha_{0}}}+o\left(t^{-\frac{2}{\alpha}}\right)\right]$
    $=$ $c_{i}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{1}{\alpha_{0}}}+c_{j}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{1}{\alpha_{0}}}+c_{i}^{\prime}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{\alpha}{\alpha_{0}}}+c_{j}^{\prime}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{\alpha}{\alpha_{0}}}+d_{i}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{2}{\alpha_{0}}}+d_{j}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{2}{\alpha_{0}}}$
    $+c_{ij}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{1}{\alpha_{0}}}-\frac{c_{i}^{2}}{2}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{2}{\alpha_{0}}}-\frac{c_{j}^{2}}{2}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{2}{\alpha_{0}}}-\frac{c_{i}^{\prime 2}}{2}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{2\alpha}{\alpha_{0}}}-\frac{c_{j}^{\prime 2}}{2}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{2\alpha}{\alpha_{0}}}$
    $-c_{i}c_{j}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{1}{\alpha_{0}}}-c_{i}c_{i}^{\prime}\frac{q_{i}^{-\frac{1+\alpha}{\alpha_{0}}}(t)}{\tau^{-\frac{1+\alpha}{\alpha_{0}}}}-c_{i}c_{j}^{\prime}\frac{q_{i}^{-\frac{1}{\alpha_{0}}}(t)q^{-\frac{\alpha}{\alpha_{0}}}_{j}(t)}{\tau^{-\frac{1+\alpha}{\alpha_{0}}}}-c_{i}^{\prime}c_{j}\frac{q_{i}^{-\frac{\alpha}{\alpha_{0}}}(t)q^{-\frac{1}{\alpha_{0}}}_{j}(t)}{\tau^{-\frac{1+\alpha}{\alpha_{0}}}}-c_{j}c_{j}^{\prime}\frac{q_{j}^{-\frac{1+\alpha}{\alpha_{0}}}(t)}{\tau^{-\frac{1+\alpha}{\alpha_{0}}}}$
    $-c_{i}^{\prime}c_{j}^{\prime}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{\alpha}{\alpha_{0}}}+o\left(t^{-\frac{2}{\alpha}}\right).$
    Then we utilize the asymptotic approximation of the marginal distribution in Eq. (A.2) to get
    $\begin{split}1-F_{ij}\{q_{i}(t),&q_{j}(t)\}=\bar{F}_{i}\{q_{i}(t)\}+\bar{F}_{j}\{q_{j}(t)\}+c_{ij}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{1}{\alpha_{0}}}\\ &-c_{i}c_{j}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{1}{\alpha_{0}}}-c_{i}c_{j}^{\prime}\frac{q_{i}^{-\frac{1}{\alpha_{0}}}(t)q^{-\frac{\alpha}{\alpha_{0}}}_{j}(t)}{\tau^{-\frac{1+\alpha}{\alpha_{0}}}}-c_{i}^{\prime}c_{j}\frac{q_{i}^{-\frac{\alpha}{\alpha_{0}}}(t)q^{-\frac{1}{\alpha_{0}}}_{j}(t)}{\tau^{-\frac{1+\alpha}{\alpha_{0}}}}\\ &-c_{i}^{\prime}c_{j}^{\prime}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{\alpha}{\alpha_{0}}}+o\left(t^{-\frac{2}{\alpha}}\right).\end{split}$
    By definition, $t^{-1}=\bar{F}_{i}\{q_{i}(t)\}=\bar{F}_{j}\{q_{j}(t)\}$. Therefore, it straightforwardly follows that $\chi_{ij}=\lim_{t\rightarrow\infty}2-t\left[1-F_{ij}\{q_{i}(t),q_{j}(t)\}\right]=0$. By Eq. (A.9) and (A.13),
@@
    $\begin{split}-\frac{1}{\eta_{ij}}=\lim_{t\rightarrow\infty}\frac{\log\left[c_{i}^{\prime}c_{j}^{\prime}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{2\alpha}{\alpha_{0}}}+o\left(t^{-2}\right)\right]}{\log t}=-2,\end{split}$
    and $\eta_{ij}=1/2$.
    If $\mathcal{C}_{i}\cap\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$,
    $\sum_{k\in\mathcal{D}}$$\left\{\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}$
    $=$ $\sum_{k\in\mathcal{D}}\left[\frac{\omega_{ki}^{1/\alpha}t^{-1/\alpha}}{c^{\prime}_{i}{}^{1/\alpha}\left\{1+R^{*}_{i}(t)+O(t^{-1/\alpha})\right\}^{1/\alpha_{0}}}+\frac{\omega_{kj}^{1/\alpha}t^{-1/\alpha}}{c^{\prime}_{j}{}^{1/\alpha}\left\{1+R^{*}_{j}(t)+O(t^{-1/\alpha})\right\}^{1/\alpha_{0}}}\right]^{\alpha}$
    $=$ $\sum_{k\in\mathcal{D}}\left(\frac{\omega_{ki}^{1/\alpha}}{c^{\prime}_{i}{}^{1/\alpha}}+\frac{\omega_{kj}^{1/\alpha}}{c^{\prime}_{j}{}^{1/\alpha}}\right)^{\alpha}t^{-1}+O(t^{1-\frac{2}{\alpha}})=d_{ij}t^{-1}+O(t^{1-\frac{2}{\alpha}}),\;\text{as }t\rightarrow\infty,$
    in which we use the asymptotic approximation in Eq. (A.13) again and by the subadditivity of power function with $\alpha\in(0,1)$,
    $d_{ij}=\sum_{k\in\mathcal{D}}\left(\frac{\omega_{ki}^{1/\alpha}}{c^{\prime}_{i}{}^{1/\alpha}}+\frac{\omega_{kj}^{1/\alpha}}{c^{\prime}_{j}{}^{1/\alpha}}\right)^{\alpha}<\sum_{k\in\mathcal{D}}\frac{\omega_{ki}}{c^{\prime}_{i}}+\sum_{k\in\mathcal{D}}\frac{\omega_{kj}}{c^{\prime}_{i}}=2.$
    Here the inequality is strict because $\mathcal{C}_{i}\cap\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$. Meanwhile $d_{ij}>\sum_{k\in\mathcal{D}}{\omega_{ki}}/{c^{\prime}_{i}}=1$. On the other hand, we note that in Eq. (A.12),
    $c_{i}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{1}{\alpha_{0}}}+c_{j}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{1}{\alpha_{0}}}=\left(\frac{c_{i}}{c^{\prime}_{i}{}^{1/\alpha}}+\frac{c_{j}}{c^{\prime}_{j}{}^{1/\alpha}}\right)t^{-\frac{1}{\alpha}}+O(t^{1-\frac{2}{\alpha}}),$
    which results in
@@
    as $t\rightarrow\infty$. Since $d_{ij}\in(1,2]$, we know from (A.8) that $\chi_{ij}=2-d_{ij}\in(0,1)$ and
@@
is a limiting measure that occurs in the limiting distribution for normalized maxima (Huser and Wadsworth, 2019). It is used to describe the multivariate extremal dependence of a spatial process, and the $n_{s}$-dimensional extremal coefficient $V(1,\ldots,1)$ is of particular interest. This extremal coefficient has a range of $[1,n_{s}]$, with the lower and upper ends indicating, respectively, perfect dependence and independence. As a polarized case, if $\gamma_{k}>0$, for all $k=1,\ldots,K$, then $\mathcal{C}_{j}\cap\mathcal{D}=\emptyset$ for all $j$’s, and thus we have
@@
Here, we can approximate $q_{j}(t)$ using the results from Corollary A.1.1. From Proposition 3.1 of the main paper, we can deduce that $V(1,\ldots,1)=n_{s}$, which corresponds to joint extremal independence. By contrast, if all $\gamma_{k}=0$ and one knot covers the entire spatial domain, we have $V(1,\ldots,1)\in[1,n_{s})$, which corresponds to joint extremal dependence.
Appendix B Validation framework details
B.1 Full range evaluation
To examine the quality of the emulation from the XVAE, we will predict at $n_{h}$ locations $\{\boldsymbol{h}_{i}:i=1,\ldots,n_{h}\}$ held out from the analyses. To perform these predictions, we calculate the basis function values at these locations, with which we can mix the encoded variables from Eq. (8) in the main paper to get predicted values. For each time $t$ and holdout location $\boldsymbol{h}_{i}$, denote the true observation of $X_{t}(\boldsymbol{h}_{i})$ by $x_{it}$ and the emulated prediction by $x^{*}_{it}$. Then the mean squared prediction error (MSPE) for time $t$ is
$\mathrm{MSPE}_{t}=\frac{1}{n_{h}}\sum_{i=1}^{n_{h}}(x_{it}-x^{*}_{it})^{2},$
where $t=1,\ldots,n_{t}$. All MSPEs from different time replicates can be summarized in a boxplot; see Section 4 in the main paper for example. Similarly, we can calculate the continuously ranked probability score (CRPS; Matheson and Winkler, 1976; Gneiting and Raftery, 2007) across time for each location, i.e.,
where $F_{i}$ is the marginal distribution estimated using parameters at the holdout location $\boldsymbol{h}_{i}$, $i=1,\ldots,n_{h}$, and again $x^{*}_{it}$ is the emulated value. Smaller CRPS indicates that the distribution $F_{i}$ is concentrated around $x^{*}_{it}$, and thus can be used to measure how well the distribution fits all emulated values. Section 4 in the main paper also shows how we present the CRPS values from all holdout locations for each emulation. In addition, we will examine the quantile-quantile (QQ)-plots obtained by pooling the spatial data into the same plot to check if the spatial input and the emulation have similar ranges and quantiles.
B.2 Empirical tail dependence measures
To assess the tail dependence structure of the emulated fields, we will estimate $\chi_{ij}(u)$ defined in Eq. (1) empirically in two ways. First, to examine the overall dependence strength, we treat $\{X(\boldsymbol{s})\}$ as if it had a stationary and isotropic dependence structure so that $\chi_{ij}(u)\equiv\chi_{h}(u)$, with $h=||\boldsymbol{s}_{i}-\boldsymbol{s}_{j}||$ being the distance between locations. Then for a fixed $h$, we find all pairs of locations with similar distances (within a small tolerance, say $\epsilon=0.001$), and compute the empirical conditional probabilities $\widehat{\chi}_{h}(u)$ at a grid of $u$ values. Confidence envelopes can be calculated by regarding the outcome (i.e., simultaneously exceed $u$ or not) of each pair as a Bernoulli variable and computing pointwise binomial confidence intervals, assuming that all pairs of points are independent from each other. Examples in Section 4 of the main paper demonstrate how this empirical measure can be used to compare the extremal dependence structures between the spatial data input and realizations from the emulator. While this metric does not completely characterize the non-stationarity in the process, it is still well-defined as a summary statistic and carries important information about the average decay of dependence with distance irrespective of the direction.
Second, to avoid the stationary assumption, we can choose a reference point denoted by $\boldsymbol{s}_{0}$ and estimate the pairwise $\chi_{0j}(u)$ empirically between $\boldsymbol{s}_{0}$ and all observed locations $\boldsymbol{s}_{j}$ in the spatial domain $\mathcal{S}$. These pairwise estimates can then be presented using a raster plot (if gridded) or a heat plot. Section 5 in the main paper shows examples of the empirical $\chi_{0j}(u)$, $u=0.85$, estimated from the real and emulated datasets, where $\boldsymbol{s}_{0}$ is the center of $\mathcal{S}$.
@@ C.1 Monte Carlo estimates of $\mathrm{ARE}_{\psi}(u)$
Proof of Theorem 3.4 of the main paper.
@@
The asymptotic normality of $n^{1/2}\{\widehat{\mathrm{ARE}}_{\psi}(u)-\mathrm{ARE}_{\psi}(u)\}$ is also ensured by Expression (C.5). However, the exact expression of its asymptotic variance requires a much more careful examination of the correlations among the ranks of $\boldsymbol{X}_{i}$, $i=0,1,\ldots,n_{g}$; that is, we need to device a multivariate linear rank statistics of $\boldsymbol{X}_{i}$, $i=0,1,\ldots,n_{g}$; see Ruymgaart and van Zuijlen (1978).
@@ C.2 Convergence of $\mathrm{ARE}_{\psi}(u)$
Proof of Theorem 3.5 of the main paper.
By the definition of the tail dependence measure in Eq. (1) of the main paper,